Terry is buying juice. He needs 3 L. A half liter of juice cost $2.39. A 250 mL container of juice c Play needs 3 L. A half liter of juice cost $2.39. A 250 mL container of juice costs $1.69.What should Terry buy so he gets 3 Liters at the lowest price? Explain
1
Expert's answer
2018-05-23T02:27:08-0400
Terry is buying juice. He needs 3 L. A half liter of juice cost $2.39.
A 250 mL container of juice costs $1.69.
One liter of juice in half liter packs costs $2.39*2 = $4.78.
One liter of juice in 250 mL packs costs $1.69*4 = $6.76.
So, Terry should buy 6 half liter packs of juice and spend
$2.39*6 = $9.56 to get 3 liters at the lowest price.